You want to create a portfolio equally as risky as the market, and you have $500,000 to invest. Information about the possiblplease help! thank you!
Answer #1

Let investment in C=$x

Hence investment in risk free asset=500,000-(137,000+143,000+x)=$(220,000-x)

Portfolio beta=Respective beta*Respective weight

1=(137,000/500,000*0.82)+(143,000/500,000*1.27)+(x/500,000*1.42)+(220,000-x)/500,000*0[Beta of market=1;Beta of risk-free assets=0]

1=0.5879+(x/500,000*1.42)

x=(1-0.5879)*500,000/1.42

=$145,105.63(Approx)=investment in C

Hence investment in risk free asset=$(220,000-x)

=$74,894.37(Approx).

Beta of risk free asset=0.